Two simple pendulums A and B are made to
oscillate simultaneously and it is found that A
completes 10 oscillations in 20 sec and B
completes 8 oscillations in 10 sec. The ratio of
the lengths of A and B will be ?

Grade:12

1 Answers

Khimraj
6 years ago
Time period in SHM is given by,
T = 2\prod \sqrt{l/g}
So l \propto T2
Time period of A, Ta = 20/10 = 2 sec.
Time period of B, Tb = 10/8 =  5/4 sec.
Ration of length La/Lb = (Ta/Tb)2 = (8/5)2 = 64/25